# Faster Cascades via Speculative Decoding

### v1

@@ C.4 Equivalence between (7) and (8)
where we choose $\alpha=0$ when $c_{2}\leq B$ and choose an $\alpha>\frac{1}{c_{2}}\cdot(c_{0}-c_{1})$. This unconstrained optimization problem is of the form in (8).
Appendix D Additional Experimental Details
We provide additional details about our experimental setup and additional experimental results.
Figure 3: Plots of quality vs. latency. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. All results use temperature sampling with $T=1$.
Figure 4: Plots of quality vs. latency. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. All results use greedy decoding ($T=0$)
Figure 5: Plots of quality vs. latency comparing lossy speculative decoding with $\beta=1$ [Lossy] against $\beta$ tuned using the procedure in [44] [Lossy].
D.1 Experimental setup and hyper-parameters
@@
Datasets. For the WMT English to German translation task [4], we use a validation sample of size 3000 provided with the dataset. We set the maximum input length to 80 and the maximum output length to 80. For the Extreme Summarization (XSum) task [31], we use a validation sample of size 11305, and set the maximum input length to 1024 and the maximum output length to 64. For the CNN/Daily Mail summarization task [20], we use a validation sample of size 13368, and set the maximum input length to 2048 and the maximum output length to 128. Following [52], we use ROUGE-2 as the evaluation metric for the summarization tasks.
@@
Models. We construct cascades from T5 v1.1 family of encoder-decoder models [34], of different sizes T5-small (77M), T5-base (250M), T5-large (800M) and T5-XL (3B). [^1] The pre-trained checkpoints we use are available at: https://console.cloud.google.com/storage/browser/t5-data/pretrained_models. We follow the protocol in [52]: we initialize with the public checkpoints, pre-train them further for 100K steps, and supervise finetune pre-trained models on the three respective tasks. We finetune them for a maximum of 250K steps on WMT, a maximum of 100K steps on XSum and a maximum of 200K steps on CNNDM.
@@
Evaluation. For each dataset, we evaluate the quality metrics on the entire validation set. For the run-time analysis, we adopt the protocol followed in [26, 52]. We randomly sample 500 examples from the validation set, and calculate the wall-clock time taken for decoding with a batch size of 1. We repeat this for three trials and report the average running time. All methods are run on the same TPUv4 device. The drafter and verifier models are run without model parallelism.
@@
Hyper-parameters. We set the block-size $\gamma$ to 5 for all methods that use speculative execution. For the token-level cascades, we allow the small model to predict for a maximum of 10 tokens (similar to [24]), before invoking the large model. This was needed, as otherwise, the small model would predict a long sequence, and when it eventually defers to the large model, the large model is bottle-necked by the pre-filling of the long prefix accumulated by the small model. We vary the lenience parameter $\alpha$ to vary the latency and plot quality as a function of latency. We vary this parameter in the range 0 to 1 for all methods where the thresholding is on a probability metric; the exceptions to this are the BiLD variants, for which, we use a longer range, as detailed below.
@@
BiLD baseline. For the BiLD method, we adopt the same discrepancy metric $D$ as [24] for greedy decoding:
$D(q,p)=\log\left(p\left(\operatorname{argmax}_{v}q(v)\right)\right),$
@@
$D(q,p)=\mathbb{E}_{v\sim q}\left[\log(p(v))\right]=\sum_{v\in\mathscr{V}}q(v)\cdot\log(p(v)).$
Lossy speculative decoding. See §D.3 for details.
D.2 Additional experimental plots
@@
In Figures 3 and 4, we provide additional plots of quality vs. latency for different inference strategies under temperature sampling ($T=1$) and greedy decoding respectively. Greedy decoding is seen to perform better compared to temperature sampling on the particular quality metrics we care about.
D.3 Lossy speculative decoding variants
@@
In §6, we compared against the lossy speculative decoding [44, 52] described in §2, with the parameter $\beta$ set to 1. We now present results for this method with $\beta$ tuned according to the procedure in [44], and show that choosing $\beta=1$ fares at least as well as tuning $\beta$.
@@
The goal in [44] is to choose $\alpha$ and $\beta$ so as to maximize the acceptance rate for the draft token, while ensuring that the KL divergence between the resulting target distribution and $p$ is within an allowable limit $R$. The authors prescribe specifying $R$, and for each prefix, tuning $\alpha$ and $\beta$ to solve the resulting constrained optimization problem. To be consistent with the rest of our experimental setup, we vary $\alpha$ to vary the draft acceptance rate (note that each choice of $\alpha$ corresponds to a particular KL divergence to $p$), and tune $\beta\geq 1-\alpha$ to satisfy the following condition outlined in [44]:
@@
In both cases, setting $\beta=1$ provides trade-offs comparable to or better than using a tuned value of $\beta$. The reason using a tuned value of $\beta$ fares worse than setting $\beta=1$ might be because we are measuring quality in terms of BLEU or ROUGE-2, which is different from the KL divergence to $p$ objective that the tuning procedure in [44] seeks to optimize.

### v2

@@ C.4 Equivalence between equation 7 and equation 8
where we choose $\alpha=0$ when $c_{2}\leq B$ and choose an $\alpha>\frac{1}{c_{2}}\cdot(c_{0}-c_{1})$ otherwise. This unconstrained optimization problem is of the form in equation 8.
Appendix D Token-specific Speculative Cascade
We provide a modification of Algorithm 5 to accommodate the token-specific deferral rules in §4.4.
Algorithm 6 TokenSpecCascade
Models $q$, $p$, Token-specific deferral rule $r$, Prefix $x_{<t}$, Block size $\gamma$
$\mathbb{T}_{\tt Token}(q,p)(v)\stackrel{{\scriptstyle.}}{{=}}q(v)\cdot(1-r(x_{<t},v))+p(v)\cdot\sum_{v^{\prime}\in\mathscr{V}}r(x_{<t},v^{\prime})\cdot q(v^{\prime})$
$\text{\tt{GenSpecSample}}(q,p,\mathbb{T}_{\tt Token},x_{<t},\gamma)$
Optimal token-specific deferral. Similar to §4.3, we may consider deriving the optimal token-specific deferral rule. We start by formulating a similar optimization objective. For a fixed prefix $x_{<t}$, this would look like:
$\min_{r}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\ell(v,\pi_{\tt Token})\big{)}\Big{]}$ (27)
[…]
where $\pi_{\tt Token}(v)\stackrel{{\scriptstyle.}}{{=}}(1-r(x_{<t},v))\cdot q_{t}(v)+\eta\cdot p_{t}(v)$ is the target distribution resulting from the choice of $r$, $\eta=\sum_{v^{\prime}\in\mathscr{V}}r(x_{<t},v^{\prime})\cdot q_{t}(v^{\prime})$ is a normalization term, and $B>0$ is a budget parameter.
However, unlike §4.3, the above constrained optimization problem does not lend itself to a simple closed-form solution. In some highly simplistic special cases, we may be able to derive a solution. For example, suppose $\ell=\ell_{\text{0-1}}$, and the mode of $q_{t}$ coincides with that of $\mathbb{P}(\cdot|x_{<t})$, i.e., $\operatorname*{arg\,max}_{v}q_{t}(v)=\operatorname*{arg\,max}_{v}\mathbb{P}(v|x_{<t})$; then the optimal token-specific rule is given by $r(x_{<t},v)=0,$ for all $v\in\mathscr{V}$.
Under more realistic cases, we may not be able to derive a solution as simple as the OPT rule in equation 10. Therefore, in our experiments, we employ the three heuristic rules in equations 1315, which are motivated by the form of the simpler Diff rule in equation 5.
Appendix E Additional Experimental Details
We provide additional details about our experimental setup and additional experimental results. We will release code and an illustrative tutorial notebook along with the final manuscript.
Figure 4: Plots of quality vs. latency for T5 models with temperature $T=1$ and block size $\gamma=5$. We include T5 plots not included in Figure 2 in the main text. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model.
Figure 5: Plots of quality vs. latency for T5 models with greedy decoding with temperature $T=0$ and block size $\gamma=5$. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model.
E.1 Experimental setup and hyper-parameters
@@
T5 datasets. For the WMT English to German translation task (Bojar et al., 2014a), we use a validation sample of size 3000 provided with the dataset. We set the maximum input length to 80 and the maximum output length to 80. For the Extreme Summarization (XSum) task (Narayan et al., 2018), we use a validation sample of size 11305, and set the maximum input length to 1024 and the maximum output length to 64. For the CNN/Daily Mail summarization task (Hermann et al., 2015), we use a validation sample of size 13368, and set the maximum input length to 2048 and the maximum output length to 128. Following (Zhou et al., 2024), we use ROUGE-2 as the evaluation metric for the summarization tasks.
We note that Kim et al. (2023) report ROUGE-L metrics for CNN/DM, which generally tend to evaluate to higher values than ROUGE-2. Furthermore, most of their experimental results are with greedy decoding ($T=0$), and hence, the ROUGE-L evaluation metrics they report in their paper tend to be higher for the same T5 models when compared to our numbers for ROUGE-2 with temperature sampling.
[…]
@@
Models. We construct cascades from T5 v1.1 family of encoder-decoder models (Raffel et al., 2020a), of different sizes T5-small (77M), T5-base (250M), T5-large (800M) and T5-XL (3B). [^1] The pre-trained checkpoints we use are available here. We follow the protocol in (Zhou et al., 2024): we initialize with the public checkpoints, pre-train them further for 100K steps, and supervise finetune pre-trained models on the three respective tasks. We finetune them for a maximum of 250K steps on WMT, a maximum of 100K steps on XSum and a maximum of 200K steps on CNNDM.
We construct the Gemma cascades from instruction-tuned decoder-only v2 models. For MBPP alone we additionally experiment with pre-trained models. We use a 2B drafter, and either a 9B verifier or a 27B verifier (Team et al., 2024).
@@
Evaluation. For each dataset, we evaluate the quality metrics on the entire validation set. For the run-time analysis, we adopt the protocol followed in Leviathan et al. (2023); Zhou et al. (2024). We randomly sample 500 examples from the validation set, and calculate the wall-clock time taken for decoding with a batch size of 1. We repeat this for three trials and report the average running time. All methods are run on the same TPUv4 device. The drafter and verifier models are run without model parallelism.
@@
Hyper-parameters. We set the block-size $\gamma$ to 5 for all methods that use speculative execution. For the token-level cascades, we allow the small model to predict for a maximum of 10 tokens (similar to (Kim et al., 2023)), before invoking the large model. This was needed, as otherwise, the small model would predict a long sequence, and when it eventually defers to the large model, the large model is bottle-necked by the pre-filling of the long prefix accumulated by the small model. We vary the lenience parameter $\alpha$ to vary the latency and plot quality as a function of latency. We vary this parameter in the range 0 to 1 for all methods where the thresholding is on a probability metric; the exceptions to this are the BiLD variants, for which, we use a longer range, as detailed below.
@@
BiLD baseline. For the BiLD method, we adopt the same discrepancy metric $D$ as (Kim et al., 2023) for greedy decoding:
$D(q,p)=-\log\left(p\left(\operatorname*{arg\,max}_{v}q(v)\right)\right),$
@@
$D(q,p)=-\mathbb{E}_{v\sim q}\left[\log(p(v))\right]=-\sum_{v\in\mathscr{V}}q(v)\cdot\log(p(v)).$
In §E.5, we present comparisons between different implementations of this method.
Lossy speculative decoding. See §E.6 for details.
E.2 Additional experimental plots
@@
In Figures 4 and 5, we provide additional plots of quality vs. latency for different inference strategies under temperature sampling ($T=1$) and greedy decoding respectively.
As noted in §C.3, with greedy decoding, the OPT deferral rule coincides with the Diff deferral rule. When temperature $T\rightarrow 0$, $D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t})=1$ whenever $\operatorname*{arg\,max}_{v}p_{t}(v)\neq\operatorname*{arg\,max}_{v}q_{t}(v)$, and is zero otherwise. In this case, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule (and $\tilde{q}_{t}$ as the drafter) is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in equation 5 as the deferral rule. In other words, for greedy decoding, the optimal deferral rules for a speculative cascade coincides with that for a sequential cascade.
Note that under greedy decoding, all methods yield better quality metrics compared to their performance under temperature sampling.
E.3 Comparing speculative deferral rules under different temperatures
In Figure 6, we present latency-quality trade-off plots for T5 cascades under temperature sampling with different temperatures. We compare lossy speculative decoding with two speculative cascade deferral rules: OPT rule in equation 5 and the Token-specific rule in equation 15. We find that the gap between OPT and the Token-specific rule diminishes as the temperature decreases.
The reason the Token-specific rule fares better than OPT is because the latter compute their deferral decisions based on which of $q_{t}(\cdot)$ and $p_{t}(\cdot)$ is more peaked; this can be a disadvantage when the sampled token is not be close the distribution mode, which is likely to happen with higher temperatures. With lower temperatures, however, the sampled token is likely to be close the distribution mode, and as a result, the advantage that the Token-specific rule has over OPT diminishes.
Figure 6: Plots of quality vs. latency for T5 models with varying temperatures. Each method interleaves T5-small with T-large. The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model.
color=RoyalBlue!35!white]add plots. Hari: done!
E.4 Comparing speculative deferral rules under different block sizes $\gamma$
In Figure 7, we present latency-quality trade-off plots for T5 cascades under different block sizes $\gamma$. In each case, we find that the proposed speculative cascading techniques outperform lossy speculative decoding across different latency values. Furthermore, higher values of $\gamma$ are seen to yield a wider range of trade-offs, with lower quality operating points shifting to the left, and better quality operating points shifting to the right. For example, with XSum, SpecDecode [Lossy] with $\gamma=3$ matches the small model’s quality at  0.64 relative latency, and matches the large model’s quality at  0.85 relative latency; with $\gamma=7$, it matches the small model’s quality at an even lower latency, but practically provides no speed-up when matching the larger model’s quality. The reason a larger block size can hurt speed-up at the higher quality regime is because it can result in frequent rollbacks, thus defeating the purpose of using speculative execution.
Figure 7: Plots of quality vs. latency for T5 models with with varying block sizes $\gamma$. Each method interleaves T5-small with T-large. The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model.
Figure 8: Top: Plots of quality vs. latency comparing BiLD with the original BiLD algorithm in Kim et al. (2023) with varying maximum draft window size $\gamma$ and fallback confidence threshold $\alpha_{f}$. Bottom: Comparison of lossy speculative decoding with $\beta=1$ [Lossy] and $\beta$ tuned using the procedure in (Tran-Thien, 2023) [Lossy].
E.5 Big Little Decoder (BiLD) variants
In §6, we compared against a version of the Big Little Decoder method (Kim et al., 2023) that applied Algorithm 4 to the target distribution $\mathbb{T}_{\textup{BiLD}}$ the authors seek to mimic (§5). We now show that this version performs similarly to the original BiLD algorithm in (Kim et al., 2023).
A key difference to the original algorithm in (Kim et al., 2023) is the use of the fallback phase, where the drafter is run until its maximum predicted probability $\max_{v}q(v)<1-\alpha_{f}$, for a threshold $\alpha_{f}\in[0,1]$ (or until a maximum block size of 10 is reached), and the use of a deterministic rollback policy where the verifier rejects a draft token whenever $D(q,p)>\alpha$. In our implementation, we adopt the speculative sampling algorithm from (Leviathan et al., 2023): we do not have the fallback policy and replace the determinisic rollback policy with the rejection sampling in Algorithm 4.
Figure 8 (top) compares the original version of BiLD with the version we use in §6. We interleave between a T5-small and T5-large model on WMT, using greedy decoding ($T=0$) for inference. As prescribed by the authors (Kim et al., 2023), we use the following discrepancy metric for greedy decoding:
$D(q,p)=\log p\left(\operatorname*{arg\,max}_{v}q(v)\right).$
We compare our implementation (BiLD), where we set the block size 5 (same as our proposed speculative cascading approaches) with the original BiLD for different choices of maximum block size $\gamma$ and different fallback thresholds $\alpha_{f}$. For both methods, we vary the threshold $\alpha$ on $D(q,p)$ to vary the latency and plot the resulting BLEU score.
A higher fallback threshold $\alpha_{f}$ results in larger draft generation windows; this gives an advantage in the low latency regime, where most of the draft tokens are accepted. As a result, BiLD [$\gamma=10,\alpha=0.9$] yields the lowest latencies, but also yields lower quality. A low fallback threshold results in very small draft generation windows, and consequently, in higher latencies. This is why BiLD [$\gamma=5,\alpha=0.1$] is the slowest but yields high quality metrics.
Our implementation BiLD is seen to perform comparable to the best parameter choices for the original BiLD algorithm in Figure 8.
Note: It is worth noting that while we view $\mathbb{T}_{\rm BiLD}$ as the target distribution that algorithm in (Kim et al., 2023) seeks to mimic, the presence of the fallback phase could mean that on some inputs a output response is generated without the verification (or rollback) phase being invoked. In such cases, the output will be a sample from the drafter even if it turns out that it contains tokens for which $D(q_{t},p_{t})>\alpha$.
E.6 Lossy speculative decoding variants
@@
In §6, we compared against the lossy speculative decoding Tran-Thien (2023); Zhou et al. (2024) described in §2, with the parameter $\beta$ set to 1. We now present results for this method with $\beta$ tuned according to the procedure in Tran-Thien (2023), and show that choosing $\beta=1$ fares at least as well as tuning $\beta$.
@@
The goal in Tran-Thien (2023) is to choose $\alpha$ and $\beta$ so as to maximize the acceptance rate for the draft token, while ensuring that the KL divergence between the resulting target distribution and $p$ is within an allowable limit $R$. The authors prescribe specifying $R$, and for each prefix, tuning $\alpha$ and $\beta$ to solve the resulting constrained optimization problem. To be consistent with the rest of our experimental setup, we vary $\alpha$ to vary the draft acceptance rate (note that each choice of $\alpha$ corresponds to a particular KL divergence to $p$), and tune $\beta\geq 1-\alpha$ to satisfy the following condition outlined in Tran-Thien (2023):
@@
In both cases, setting $\beta=1$ provides trade-offs comparable to or better than using a tuned value of $\beta$. The reason using a tuned value of $\beta$ fares worse than setting $\beta=1$ might be because we are measuring quality in terms of BLEU or ROUGE-2, which is different from the KL divergence to $p$ objective that the tuning procedure in Tran-Thien (2023) seeks to optimize.
E.7 Token-specific speculative cascades
In Figure 9, we present latency-quality trade-off plots for cascades constructed from a T5 small and a T5 large model. We include in these comparisons, all three token-specific deferral rules in equations 1315. In Figure 10, we present trade-off plots for cascades constructed from Gemma 2B and Gemma 27B models with all three token-specific rules, and in Figure 11, we include similar plots for cascades constructed from Gemma 2B and Gemma 9B models. We note that the trends with the 2B $\rightarrow$ 9B are similar to those seen with the 2B $\rightarrow$ 27B cascades.
[…]
The reason the V3 variant outperforms V1 and V2 on the Gemma models could be due to the fact that it uses the larger model’s distribution $p_{t}(\cdot)$ to measure confidence for both the drafter and verifier (see LHS and RHS in equation 13). We expect this to be particularly helpful when there is a larger gap in sizes between $q$ and $p$, and the larger model’s distribution is better aligned with the ground-truth distribution compared to the smaller model. Furthermore, as noted in §4.4, the multiplicative form of the rule results in a target distribution that has an intuitive form: it seeks to mimic $q_{t}(\cdot)$ on the top-$\alpha$ ranked tokens by $p_{t}(\cdot)$ and uses a re-scaled version of $p_{t}(\cdot)$ for the other tokens.
Figure 9: Plots of quality vs. latency for T5 models with all three token-specific speculative cascade deferral rules in equations 1315. Each method interleaves a T5 small and a T5 large model. The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model.
Figure 10: Plots of quality vs. rejection rate for Gemma models with all three token-specific speculative cascade deferral rules in equations 1315. Each method interleaves a Gemma 2B drafter with a Gemma 27B verifier. The horizontal dotted line denotes the quality of the large model. We include all three token-specific speculative cascade deferral rules in equations 1315.
color=RoyalBlue!35!white]add plots. Hari:done!
Figure 11: Plots of quality vs. rejection rate with Gemma 2B $\rightarrow$ 9B speculative cascades. Each method interleaves a Gemma 2B drafter with a Gemma 9B verifier. The horizontal dotted line denotes the quality of the large model. We include all three token-specific speculative cascade deferral rules in equations 1315.
color=RoyalBlue!35!white]add plots. Hari:done!
Appendix F Limitations
One of the limitations of our proposal is the use of plug-in estimators to approximate the optimal rule equation 9. While these approximations are effective in practice, they rely on the individual models being calibrated. An alternative to the use of plug-in estimators is to use a router model explicitly trained to mimic the optimal rule using a validation sample drawn from $\mathbb{P}$ (Gupta et al., 2024). Another limitation is that the optimization objectives we seek to minimize are local objectives that seek to make the best deferral decision at the current position $t$. In doing so, they ignore the downstream effects of choosing a particular model in the current step. Devising a global deferral objective that takes downstream errors into account would be an interesting direction for future work. More broadly, our paper seeks to improve cost-quality trade-offs in LM inference. It is important that such improvements do not unfairly advantage one slice of the data or a subset of the population, at the cost of others. Ensuring that the trade-off gains that our approach offers is equitable across different slices of the data is another important direction for the future.